OpenStudy (lacedupcountry):

I have a Consumer Math/Algebra Question, if anyone can help me I'll fan and medal! I'll post it below

OpenStudy (lacedupcountry):

This is what I have so far, but I'm not sure if i am doing it right.

OpenStudy (lacedupcountry):

@amistre64 @ganeshie8

OpenStudy (amistre64):

part 2 feels off to me.

OpenStudy (amistre64):

it looks like you used an APY when calculating a formula that uses that APR

OpenStudy (amistre64):

we would expect that at a given apy and apr 500(1+apy) = 500(1+apr/n)^n or simply: 1 + apy = (1+apr/n)^n the apr has to do with the compounding interest rate; while the apy is an expected yeild

OpenStudy (amistre64):

part 3 doesnt seem right either. the function should represent total number of payments for the year; +50 or, monthly payment, plus 50/12

OpenStudy (lacedupcountry):

Okay so for question 2 i need to use that formula? 1 + apy = (1+apr/n)^n

OpenStudy (lacedupcountry):

@amistre64

OpenStudy (amistre64):

i cant say for sure what 2 needs, ive never been good at wordy questions. but as is i think your approach is off and that we would need to convert the apy to apr in order to adress it correctly

OpenStudy (amistre64):

how often the account calculates itself would be useful to determine the value of n. if its quarterly n=4, it is monthly n=12, if its twice a year n=2 etc ....

OpenStudy (amistre64):

but yes, finding the apr: 1 + apy = (1+apr/n)^n nrt (1+apy) = 1+apr/n nrt (1+apy) - 1 = apr/n n [nrt (1+apy) - 1] = apr

OpenStudy (lacedupcountry):

Thanks for your help